# DITING: A Static Analyzer for Identifying Bad Partitioning Issues in TEE Applications

### v1

@@ Abstract.
To address this, we propose an approach to detect partitioning issues in TEE applications. First, we conducted a survey of TEE vulnerabilities caused by bad partitioning and found that the parameters exchanged between the secure and normal worlds often contain insecure usage with bad partitioning implementation. Second, we developed a tool named DITING that can analyze data-flows of these parameters and identify their violations of security rules we defined to find bad partitioning issues. Different from existing research that only focuses on malicious input to TEE, we assess the partitioning issues more comprehensively through input/output and shared memory. Finally, we created the first benchmark targeting bad partitioning, consisting of 110 test cases. Experiments demonstrate the DITING achieves an F1 score of 0.90 in identifying bad partitioning issues.
@@
Trusted Execution Environment, bad partitioning, static analysis, data-flow, shared memory
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Software and application securityccs: Security and privacy Software security engineeringccs: Software and its engineering Software evolution
1. Introduction
@@
Trusted Execution Environment (TEE) has emerged as an essential component in enhancing the security of mobile applications and cloud services (Huang et al., 2024; Santos et al., 2014; Dong et al., 2023). In a typical TEE application, code is usually divided into secure and non-secure parts. The non-secure code operates in the non-secure environment, also known as the normal world, while the secure code runs within TEE. This mechanism provides significant security advantages by isolating the secure code from the normal world, preventing attackers from accessing sensitive data or disrupting secure functions, such as key management (Luo et al., 2018; Park et al., 2020), cryptography (Volante et al., 2024; Vella et al., 2021), and memory operations (Zhao and Mannan, 2019; Feng et al., 2024). To achieve this isolation, TEE leverages hardware-based features like memory partitioning and access control enforced by the CPU (Zhang et al., 2024; Atamli-Reineh et al., 2017). When the code in the normal world needs to access secure functions or data, it must interact with TEE using specific interfaces. Therefore, partitioning strategies of TEE applications that incorporate accurate isolation (Zhao et al., 2019; Geneiatakis et al., 2012; Kim et al., 2018) and access control (Liu et al., 2023; Geneiatakis et al., 2012) mechanisms can be helpful in reducing the attack surface. For instance, Zhao et al. (Zhao et al., 2019) designed SecTEE, a solution that integrates computing primitives (e.g., integrity measurement, remote attestation, and life cycle management) into TEE. This design enables SecTEE to resist privileged host software attacks from the normal world through isolation. Liu et al. (Liu et al., 2023) proposed a multi-owner access control scheme for access authorization in Intel SGX, which can prevent unauthorized data disclosure. The above research has shown that well-designed partitions of TEE applications can help mitigate risks by establishing clear boundaries and secure communication protocols between the normal world and TEE.
@@
However, TEE applications still face some threats. For example, Khandaker et al. (Khandaker et al., 2020) summarized four attacks that can be implemented in Intel SGX: Concurrency, Order, Inputs, and Nested (COIN attacks). In these attacks, malicious TEE users can exploit the lax detection vulnerabilities of non-secure inputs (e.g., buffer and value with incorrect size) in Intel SGX to interfere with executing programs in TEE. Cerdeira et al. (Cerdeira et al., 2020) conducted a security investigation on some popular systems based on TrustZone. They found that input validation weaknesses and other bugs in TrustZone-assisted systems can be used to obtain permissions of TEE. For example, Huawei TEE allows a TEE application to dump its stack trace to the normal-world memory through debugging channels, leaking sensitive information.
Figure 1. A bad partitioning case of input validation weaknesses.
@@
Through analysis of these cases, we found that vulnerabilities in TEE applications often arise from insecure data communication between the normal world and TEE (Fleischer et al., 2020; Jang and Kang, 2020), which usually relies on three types of parameters: input/output and shared memory (Fan et al., 2020). These parameters are typically managed by the normal world, making them susceptible to unauthorized access. Hence, the usage of these parameters without isolation and access control (e.g., unencrypted data output, input validation weaknesses, and direct usage of shared memory) can introduce bad partitioning issues to TEE applications. For example, there is a bad partitioning case in Fig. 1. The normal-world code validates the size of the input buffer before invoking TEE, but this validation step is missing inside the TEE code. While attackers who gain control over the normal world have the ability to intercept sensitive data or gain unauthorized access to secure operations (Cerdeira et al., 2022), they can tamper with the size of the input buffer before the TEE invocation and disrupt the execution of the TEE application through malicious input. Such scenarios highlight the critical importance of ensuring that partition boundaries are well-defined and communication channels are fortified against potential breaches. Therefore, this paper discusses TEE bad partitioning issues based on two key aspects: improperly ported or unported code into TEE, and the lack of in-TEE code which securely interfaces the normal world and TEE. Unfortunately, there are currently no tools to analyze and identify TEE partitioning issues.
@@
To address these challenges and enhance TEE security, we formulated three rules for detecting these issues. Following these rules, we designed a data-flow analyzer to identify bad partitioning in TEE applications by determining whether input/output and shared memory parameters passed between the normal world and TEE violate the defined rules. This is the first automatic tool that can identify bad partitioning issues in TEE applications. Our contributions of this paper include:
@@
  • The first systematic summary of partitioning issues. With the real-world vulnerable code where input, output, and shared memory parameters are handling in sensitive functions and data of TEE, we summarized that the bad partitioning issues in TEE are often caused by the improper usage of these parameters. Unlike previous research (Khandaker et al., 2020; Cerdeira et al., 2020), which primarily focused on malicious inputs to TEE, our approach addresses the potential vulnerabilities brought by three parameters within TEE, offering a more comprehensive solution to bad partitioning issues.
@@
  • Proposing a parameter data-flow analyzer to identify partitioning issues. Based on parameter types, we established a set of rules that these parameters must adhere to as they flow through sensitive functions and and interact with sensitive data. By applying data-flow analysis to determine whether these parameters violate the established rules caused by partitioning issues, we can effectively identify instances of bad partitioning in TEE applications. We named it DITING, a fantastic beast has the ability to distinguish between good and evil from Chinese Mythology Journey to the West.
@@
  • Creating the benchmark for bad partitioning detection. Due to the lacking of datasets containing TEE partitioning issues, we created the first Partitioning Errors Benchmark named PartitioningE-Bench, which consists of 110 test cases, and 90 cases of them cover 3 types of bad partitioning vulnerabilities. Experimental results demonstrate that DITING achieves an F1 score of 0.90 in identifying bad partitioning issues. Using DITING, we also find 55 issues in the real-world programs and are confirming them with the developers.
  • The source code of DITING and PartitioningE-Bench are available on https://github.com/CharlieMCY/PartitioningE-in-TEE.
@@
The rest of the paper is organized as follows. We first describe the data interaction model and bad partitioning issues of TEE applications in Section 2. Next, Section 3 presents the system design of DITING and Section 4 evaluates the system. Subsequently, we discuss the threats to validity and related works in Section 5 and Section 6, respectively. Finally, we conclude the paper and mention future work in Section 7.
2. Background and Motivation
2.1. TEE Data Interaction
@@
As shown in Fig. 2, the normal world and TEE are two independent environments, separated to ensure the security of sensitive functions and data. In this architecture, the communication between TEE and the normal world involves three types of parameters: input, output, and shared memory (Fan et al., 2020).
Figure 2. Data communication between TEE and the normal world.
@@
Shared memory provides a zero-copy memory block to exchange larger data sets (e.g., multimedia files or bulk data) between two sides (TrustedFirmware.org, 2025). It allows both sides to access the same memory space efficiently, which avoids frequent memory copying. Shared memory can also remain valid across different TEE invocation sessions, making it suitable for scenarios that require data to be reused multiple times.
@@
Moreover, an SDK is responsible for managing these parameters and communication in the normal side. For example, TrustZone-based OP-TEE uses TEEC_InvokeCommand() function to execute the in-TEE code, enabling the shared memory or temporary buffers to transfer data between the two sides (Sebastian et al., 2019). Then, OP-TEE can handle these interactions through its internal APIs, which process incoming requests, perform secure computations, and return responses to the normal world. Similarly, Intel SGX utilizes the ECALL and OCALL interfaces to achieve these functionalities (Regano and Canavese, 2024). ECALLs allow the normal world to securely invoke functions within the SGX enclave, passing data into the trusted environment for processing, while OCALLs enable the enclave to request services or share results with the untrusted normal world.
Table 1 illustrates some differences between input/output parameters and shared memory. It is important to note that, while the normal side cannot directly access the memory copies of input and output parameters in TEE, an attacker with the permissions of the normal world can still tamper with the inputs before they are transmitted to TEE or intercept and read the outputs after they are returned from TEE (Lu et al., 2023; Ma et al., 2024). Additionally, since shared memory relies on address-based data transfer between the two sides, any modifications made to the data on one side will be instantly mirrored on the other opposing side. Therefore, data interactions controlled by the normal world code are the root cause of the bad partitioning issues in TEE applications.
Table 1. Comparison of input/output parameters and shared memory.
Feature Input/Output Shared Memory
Data Size Small Big
Efficiency Low, memory copying required High, no need of additional memory copy
Life-time Temporary, valid for a single TEE invocation Valid for a long time and shared among several TEE invocations
Privilege The normal world cannot reach memory copy in TEE Both the normal world and TEE can access at the same time
2.2. Bad Partitioning in TEE Projects
By analyzing 10 TEE projects from github.com, we identified three primary types of bad partitioning issues: ① unencrypted data output, ② input validation weaknesses, and ③ the direct usage of shared memory. We list some projects and issues they face in Table 2, and discuss each of these issues in detail below. In this paper, we use projects developed for ARM TrustZone as examples. The issues we identify and the proposed solutions work for other platforms as well.
@@ 2.2. Bad Partitioning in TEE Projects
Table 2. Surveys of bad partitioning issues in real-world TEE projects. ● means that the project contains the corresponding issue, and ○ indicates not containing the issue.
@@
2.2.1. Unencrypted Data Output
@@
(b) TEE data delta_refcount is directly copied to the output buffer.
Figure 3. The example of unencrypted data output from optee-sdp.
1char g_AesOutpUT[256] = {0};
2// op.params[1].tmpref.buffer is an input buffer larger than test buffer in TEE
@@
4res = l_CryptoVerifyCa_SendCommand(&op, &session, CMD_AES_OPER);
(a) The input buffer and its size are defined outside TEE.
1char test[] = {0x01, 0x02, 0x03, 0x04, 0x05};
2// params[1].memref.buffer is the copy of op.params[1].tmpref.buffer in the normal world
3TEE_MemMove(params[1].memref.buffer, test, sizeof(test));
(b) TEE_MemMove is called without verifying the size of input buffer.
Figure 4. The example of input validation weaknesses from basicAlg_use.
@@
1// params[2].memref.buffer is shared memory
@@
4// params[2].memref.buffer may have been changed outside TEE
@@
(b) TEE_MemCompare is invoked without checking the buffer contents.
@@
Figure 5. The example of the direct usage of shared memory from external_rk_tee_user.
@@
Fig. 3 shows a scenario where if the TEE data is copied to an output buffer without being encrypted, then the attacker can directly obtain the plaintext from dump in the normal world. Therefore, implementing proper security measures is essential to prevent unauthorized access or data leakage during the TEE output.
2.2.2. Input Validation Weaknesses
@@
As mentioned in Section 2.1, an attacker with normal-world permissions has the ability to tamper with any data intended for input into TEE. Fig. 4 illustrates that even though the normal-world code defines a much larger buffer than text buffer used in TEE, it is still susceptible to buffer overflow attacks that in-TEE code directly performs memory copy operations in the TEE code without validating the size of the input buffer params[1].memref.buffer. This occurs because the TEE code may incorrectly assume that the input buffer is secure and reliable without explicitly checking its size. If an attacker maliciously alters the input buffer size, memory copy operations within TEE could exceed the boundaries of the inputs, leading to severe consequences such as crashes and data corruption. Similar precautions should be taken when using input values as indices for array access in TEE. Moreover, checking the input value treated as the size argument of the memory request function TEE_Malloc(size) can effectively reduce the probability of memory allocation failure.
To mitigate this risk, it is necessary for the TEE code to validate all inputs before performing any memory operations, regardless of whether input checking is partitioned to the normal side. This essential validation ensures that any tampering or unexpected changes to the input data in the normal world do not compromise the security of TEE.
2.2.3. Direct Usage of Shared Memory

### v2

@@ Abstract
To address this, we propose an approach to detect partitioning issues in TEE applications. First, we conducted a survey of TEE vulnerabilities caused by bad partitioning and found that the parameters exchanged between the secure and normal worlds often contain insecure usage with bad partitioning implementation. Second, we developed a tool named DITING that can analyze data flows of these parameters and identify their violations of security rules we defined to find bad partitioning issues. Different from existing research that only focuses on malicious input to TEE, we assess the partitioning issues more comprehensively through input/output and shared memory. Finally, we created the first benchmark targeting bad partitioning, consisting of 110 test cases. Experiments demonstrate that DITING achieves an F1 score of 0.90 in identifying bad partitioning issues.
Index Terms:
@@
Trusted Execution Environment, bad partitioning, static analysis, data flow, shared memory.
I Introduction
@@
Trusted Execution Environment (TEE) has emerged as an essential component in enhancing the security of mobile applications and cloud services [1, 2, 3]. In a typical TEE application, code is usually divided into secure and non-secure parts. The non-secure code operates in the non-secure environment, also known as the normal world, while the secure code runs within a TEE. This mechanism provides significant security advantages by isolating the secure code from the normal world, preventing attackers from accessing sensitive data or disrupting secure functions, such as key management [4, 5], cryptography [6, 7], and memory operations [8, 9]. To achieve this isolation, TEE leverages hardware-based features like memory partitioning and access control enforced by the CPU [10, 11]. When the code in the normal world needs to access secure functions or data, it must interact with TEE using specific interfaces. Therefore, partitioning strategies of TEE applications that incorporate accurate isolation [12, 13, 14] and access control [15, 13] mechanisms can be helpful in reducing the attack surface. For instance, Zhao et al. [12] designed SecTEE, a solution that integrates computing primitives (e.g., integrity measurement, remote attestation, and life cycle management) into TEE. This design enables SecTEE to resist privileged host software attacks from the normal world through isolation. Liu et al. [15] proposed a multi-owner access control scheme for access authorization in Intel SGX, which can prevent unauthorized data disclosure. The above research has shown that well-designed partitions of TEE applications can help mitigate risks by establishing clear boundaries and secure communication protocols between the normal world and TEE.
@@
However, TEE applications still face some threats. For example, Khandaker et al. [16] summarized four attacks that can be implemented in Intel SGX: Concurrency, Order, Inputs, and Nested (COIN attacks). In these attacks, malicious TEE users can exploit the lax detection vulnerabilities of non-secure inputs (e.g., buffer and value with incorrect size) in Intel SGX to interfere with executing programs in TEE. Cerdeira et al. [17] conducted a security investigation on some popular systems based on TrustZone. They found that input validation weaknesses and other bugs in TrustZone-assisted systems can be used to obtain permissions of TEE. For example, Huawei TEE allows a TEE application to dump its stack trace to the normal-world memory through debugging channels, leaking sensitive information.
Figure 1: A bad partitioning case of input validation weaknesses.
@@
Through analysis of these cases, we found that vulnerabilities in TEE applications often arise from insecure data communication between the normal world and TEE [18, 19], which usually relies on three types of parameters: input/output and shared memory [20]. These parameters are typically managed by the normal world, making them susceptible to unauthorized access. Hence, the usage of these parameters without isolation and access control (e.g., cryptographic functions and bounds checking) can introduce bad partitioning issues to TEE applications. For example, there is a bad partitioning case in Fig. 1. The normal-world code validates the size of the input buffer before invoking TEE, but this validation step is missing inside the TEE code. If an attacker controls the normal world and has the ability to intercept sensitive data or gain unauthorized access to secure operations [21], they can tamper with the size of the input buffer before the TEE invocation. The TEE application is then vulnerable to buffer overflow attacks caused by malicious inputs. Such scenarios highlight the critical importance of ensuring that partition boundaries are well-defined and communication channels are fortified against potential breaches. Therefore, this paper discusses TEE bad partitioning issues based on two key aspects: (1) the absence of secure interface components between the normal world and the TEE (e.g., returning output data without encryption, and using input data missing bounds checking), and (2) the use of insecure mechanisms within TEE code (e.g., insecure reliance on shared memory). Unfortunately, there are currently no tools to analyze and identify TEE partitioning issues.
@@
To address these challenges and enhance TEE security, we formulated three rules for detecting these issues. Following these rules, we designed a data flow analyzer to identify bad partitioning in TEE applications by determining whether input/output and shared memory parameters passed between the normal world and TEE violate the defined rules. This is the first automatic tool that can identify bad partitioning issues in TEE applications. Our contributions of this paper include:
@@
  • The first systematic summary of partitioning issues. With the real-world vulnerable code where input, output, and shared memory parameters are handling in sensitive functions and data of TEE, we summarized that the bad partitioning issues in TEE are often caused by the improper usage of these parameters. Unlike previous research [16, 17], which primarily focused on malicious inputs to TEE, our approach addresses the potential vulnerabilities brought by three parameters within TEE, offering a more comprehensive solution to bad partitioning issues.
@@
  • Proposing a parameter data flow analyzer to identify partitioning issues. Based on parameter types, we established a set of rules that these parameters must adhere to as they flow through sensitive functions and interact with sensitive data. By applying data flow analysis to determine whether these parameters violate the established rules caused by partitioning issues, we can effectively identify instances of bad partitioning in TEE applications. We named it DITING, a fantastic beast that has the ability to distinguish between good and evil from the Chinese Mythology novel Journey to the West.
@@
  • Creating the benchmark for bad partitioning detection. Due to the lack of datasets containing TEE partitioning issues, we created the first Partitioning Errors Benchmark named PartitioningE-Bench, which consists of 110 test cases, and 90 cases of them cover 3 types of bad partitioning vulnerabilities. Experimental results demonstrate that DITING achieves an F1 score of 0.90 in identifying bad partitioning issues. Using DITING, we also find 130 issues in 16 real-world TEE projects. Then, we have contributed some pull requests addressing the issues in 5 projects and one of them has already been confirmed and merged into the main branch by the maintainers.
@@
The rest of the paper is organized as follows. We first describe the data interaction model and bad partitioning issues of TEE applications in Section II. Next, Section III presents the system design of DITING and Section IV evaluates the system. Subsequently, we discuss the threats to validity and related works in Section V and Section VI, respectively. Finally, we conclude the paper and mention future work in Section VII.
II Background and Motivation
II-A TEE Data Interaction
@@
As shown in Fig. 2, the normal world and TEE are two independent environments, separated to ensure the security of sensitive functions and data. In this architecture, the communication between TEE and the normal world involves three types of parameters: input, output, and shared memory [20].
Figure 2: Data communication between TEE and the normal world.
@@
Shared memory provides a zero-copy memory block to exchange larger data sets (e.g., multimedia files or bulk data) between two sides [22]. It allows both sides to access the same memory space efficiently, which avoids frequent memory copying. Shared memory can also remain valid across different TEE invocation sessions, making it suitable for scenarios that require data to be reused multiple times.
@@
Moreover, an SDK is responsible for managing these parameters and communication in the normal side. For example, TrustZone-based OP-TEE uses TEEC_InvokeCommand() function to execute the in-TEE code, enabling the shared memory or temporary buffers to transfer data between the two sides [23]. Then, OP-TEE can handle these interactions through its internal APIs, which process incoming requests, perform secure computations, and return responses to the normal world. Similarly, Intel SGX utilizes the ECALL and OCALL interfaces to achieve these functionalities [24]. ECALLs allow the normal world to securely invoke functions within the SGX enclave, passing data into the trusted environment for processing, while OCALLs enable the enclave to request services or share results with the untrusted normal world.
II-B Threat Model
In this paper, we consider an adversary who has full control over the normal-world code and can tamper with the parameters passed between the normal world and TEE. This includes the ability to manipulate input values, intercept or observe output data, and modify shared memory contents. We identify three main threat vectors that emerge from bad partitioning practices:
  • Sensitive Data Leakage: If sensitive data (e.g., cryptographic keys, passwords, or authentication tokens) is output to the normal world without encryption, it may be exposed to adversaries. This risk is further exacerbated by inference attacks [25], where attackers infer protected information by observing output patterns or variations when they are returned in plain text.
  • Insecure Input Handling: The TEE SDK typically performs basic validation checks when copying parameters into TEE, such as preventing null pointers. However, the SDK lacks semantic knowledge of how those inputs are used within the TEE application [16]. Thus, the burden of input validation within TEEs, such as checking buffer sizes or ensuring value ranges, falls on the developer. If such validation is missing or incomplete, the TEE becomes vulnerable to memory security issues such as buffer overflows, even when inputs were superficially checked in the normal world.
  • Shared Memory Manipulation: Both the normal side and TEE can access the shared memory. However, existing TEEs generally lack built-in mechanisms to ensure the integrity of shared memory contents [26], and most hardware vendors do not provide standardized guidelines for secure shared memory usage [18]. As a result, a common pitfall is that TEE applications may access shared memory directly through shallow copies, making them vulnerable to runtime manipulation by untrusted normal-world processes controlled by attackers. This risk has been documented in prior studies like EnclaveFuzz [27] and prompted ongoing community discussions on best practices for securely handling shared memory in TEE systems [^1] https://github.com/lsds/sgx-lkl/issues/745.
The above vulnerabilities have also been exemplified in real-world TEE applications, such as CVE-2016-10237222https://www.cve.org/CVERecord?id=CVE-2016-10237, CVE-2020-8936333https://www.cve.org/CVERecord?id=CVE-2020-8936 and CVE-2021-0186444https://www.cve.org/CVERecord?id=CVE-2021-0186. Therefore, our threat model focuses on attacks stemming from the misuse of input/output and shared memory parameters in the TEE code, while the misuse refers to the vulnerable code that misses secure components or directly calls insecure components when using the above parameters.
@@ II-B Threat Model
TABLE I: Surveys of bad partitioning issues in real-world TEE projects. ● means that we found the corresponding issue in this project, and ○ indicates that we did not find.
@@
II-C Bad Partitioning in TEE Projects
We searched for TEE-related projects on GitHub using API keywords associated with TEE. After excluding TEE SDK projects, we selected the top 10 projects with the highest star counts. By manually analyzing these TEE projects, we identified three primary types of bad partitioning issues: ① unencrypted data output, ② input validation weaknesses, and ③ the direct usage of shared memory. We list some projects and issues they face in Table I, and discuss each of these issues in detail below. In this paper, we use projects developed for ARM TrustZone as examples. The issues we identify and the proposed solutions work for other platforms as well.
II-C1 Unencrypted Data Output
@@
(b) Data delta_refcount is directly copied to the output buffer in TEE code.
Figure 3: The example of unencrypted data output from optee-sdp.
1char g_AesOutpUT[] = {0x01, 0x02, 0x03, 0x04, 0x05};
2// op.params[1].tmpref is an input buffer smaller than test buffer in TEE
@@
4op.params[1].tmpref.size = 5;
5res = l_CryptoVerifyCa_SendCommand(&op, &session, CMD_AES_OPER);
(a) The input buffer and its size are defined in normal-world Code.
1char test[256] = {};
2// params[1].memref is the copy of op.params[1].tmpref in the normal world
3TEE_MemMove(test, params[1].memref.buffer, params[1].memref.size);
(b) TEE_MemMove is called without verifying the size of input buffer in TEE code.
Figure 4: The example of input validation weaknesses from basicAlg_use.
@@
1// params[2].memref is shared memory
@@
4// params[2].memref may have been changed outside TEE
@@
(b) TEE_MemCompare is invoked with shared memory in TEE code.
@@
Figure 5: The example of the direct usage of shared memory from external_rk_tee_user.
@@
Fig. 3 shows a scenario where if the TEE data is copied to an output buffer without being encrypted, then the attacker can directly obtain the plaintext from dump in the normal world. Therefore, implementing proper security components, such as cryptographic functions for encryption, is essential to prevent unauthorized access or data leakage during the TEE output.
II-C2 Input Validation Weaknesses
@@
As mentioned in Section II-B, an attacker with normal-world permissions has the ability to tamper with any data intended for input into TEE. Fig. 4 illustrates that even though the normal-world code defines a smaller buffer than text buffer used in TEE, it is still susceptible to buffer overflow attacks that in-TEE code directly performs memory copy operations in the TEE code without validating the size of the input buffer params[1].memref.buffer. This occurs because the TEE code may incorrectly assume that the input buffer is secure and reliable without explicitly checking its size. If an attacker maliciously alters the input buffer size, memory copy operations within TEE could exceed the boundaries of the inputs, leading to severe consequences such as crashes and data corruption. Similar precautions should be taken when using input values as indices for array access in TEE. We also call the above precautions security components, such as bounds checking for input parameters.
II-C3 Direct Usage of Shared Memory